a+b+c=12
a=5b
a+c=9

Answer by edjones(8007)

a+b+c=12
a=5b
a+c=9
.
c=9-a
c=9-5b since a=5b
.
5b+b+9-5b=12
b+9=12
b=3
a=15
c=-6
.
Ed